OpenStudy (anonymous):

Find the distance between the points (-4, -5) and (3, -1)

OpenStudy (anonymous):

use distance formula. \[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

you have x1=-4 x2=3 y1=-5 y2=-1

OpenStudy (anonymous):

put vslues and solve..to get rresult...

OpenStudy (anonymous):

My choice are A)√65 b)√37 d)√17

OpenStudy (anonymous):

Make it in to a vector AB by subtracting the distances from eachother (-4 -3 =7, =-5 - -1 =-4) or (7,4) as a vector. Now to find the length of a vector, take the square root of the absolute values. \[\sqrt{7^{2}+(-4)^{2}} = \sqrt{49 + 16} = \sqrt{65}\]

OpenStudy (anonymous):

I of course meant subtracting the points from eachother

OpenStudy (anonymous):

Awe Thanks I Get it a little